OpenStudy (anonymous):

I need help figuring out how to get the perpendicular line to -2x+3y=9 through (-1,-3)

OpenStudy (luigi0210):

First put the equation in slope-intercept form

OpenStudy (luigi0210):

After that, obtain the slope. And perpendicular means the negative reciprocal, so plug that into the point-slope formula, and if needed simplify.

OpenStudy (luigi0210):

Also\[\large \mathbb{\color{red}{WELCOME~TO~OPENSTUDY}}\]

OpenStudy (anonymous):

So the slope intercept form would be y= 2/3x +3 ? But I don't know exactly what to do after, I need a more clear explaination please.

OpenStudy (luigi0210):

Do you know the slope?

OpenStudy (anonymous):

That's what I need help on trying to figure out.

OpenStudy (luigi0210):

The slope in y=mx+b is m SO in y=(2/3)x+3 would be 2/3

OpenStudy (anonymous):

So when I put it in reciprocal form it will be -2/3 ? Right?

OpenStudy (luigi0210):

negative reciprocal: \[\Large \frac{a}{b}=-\frac{b}{a}\]

OpenStudy (anonymous):

ohh , now I understand it better. So it would be -3/2 , then I plug in the points that I have , then use the method to find b so y=-3/2x+b? or do I just put b as 0?

OpenStudy (anonymous):

Or do I do it the other way around

OpenStudy (luigi0210):

No you're right, -3/2, plug in and find in slope intercept form :)

OpenStudy (anonymous):

It would be \[y=\frac{- 3 }{2 }x+\frac{ 5 }{ 2 }\] then right?

OpenStudy (luigi0210):

Hmm, I got \[y=-\frac{3}{2}x-\frac{9}{2}\] Maybe I messed up somewhere

OpenStudy (luigi0210):

I did mess up, whoops.. but now I got -3/2 instead .-.

OpenStudy (luigi0210):

I'll show you my process if you want?

OpenStudy (anonymous):

Yea that will be kinda better because I think I did something wrong.

OpenStudy (luigi0210):

Okay, so we got the formula: \[\LARGE y-y_{1}=m(x-x_{1})\] and the points \[\Large (-1, -3)==>(x_{1}, y_{1})\] So we plug that in: \[\huge y+3=-\frac{3}{2}(x+1)\] Now distribute and get: \[\huge y+3=-\frac{3}{2}x-\frac{3}{2}\] Now subtract 3 from both sides \[\huge y+3-3=-\frac{3}{2}x -\frac{3}{2}-\frac{6}{2}\] Since 3=6/2 in fraction form, now simplify: \[\huge y=-\frac{3}{2}x-\frac{9}{2}\] Try to see where I messed up >.<

OpenStudy (anonymous):

You actually got it right, I had got it wrong because I didnt use the right formula ._.

OpenStudy (luigi0210):

Oh, well then :P

OpenStudy (luigi0210):

Glad I could help ^-^

OpenStudy (anonymous):

Thank you (:

OpenStudy (luigi0210):

Anytime :) and not trying to be selfish or anything but can I have a medal? :P

OpenStudy (luigi0210):

Thank you :D

OpenStudy (anonymous):

No Problem